# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v2

@@ Abstract
With over 700 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models. Despite its widespread use, its ability to maintain model accuracy during optimization has not been thoroughly investigated. In this work, we present OODTE, a utility designed to automatically and comprehensively evaluate the correctness of the ONNX Optimizer. OODTE adopts a straightforward yet powerful differential testing and evaluation methodology, which can be readily adapted for use with other compiler optimizers. Specifically, OODTE takes a collection of ONNX models, applies optimizations, and executes both the original and optimized versions across a user-defined input set, automatically capturing any issues encountered during optimization. When discrepancies in accuracy arise, OODTE iteratively isolates the responsible optimization pass by repeating the process at a finer granularity. We applied OODTE to 130 well-known models from the official ONNX Model Hub, spanning diverse tasks including classification, object detection, semantic segmentation, text summarization, question answering, and sentiment analysis. Our evaluation revealed that $9.2$% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies. Additionally, $30$% of classification models and $16.6$% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization. In total, OODTE uncovered 15 issues—14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly. All findings were reported to the ONNX Optimizer development team. OODTE demonstrates a simple but effective framework for validating AI model optimizers and with minimal work, is applicable beyond the ONNX ecosystem.
I Introduction
@@
Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models [13, 17]. Widely-used AI compilers and frameworks, such as Apache TVM [12] and MLIR [32], incorporate their own native sets of optimization passes for efficient use of hardware. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
@@
Models can also be represented using industry-standard formats, such as ONNX [8]. By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer [4] serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question.
@@
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. Ideally, such transformations should not introduce any deviations in model behavior. To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a framework for performing differential analysis between original and optimized models. Using OODTE, we collected 130 models of various types, including classification, object detection, text comprehension and generation from the official ONNX Model Hub [7]. Each model was optimized using ONNX Optimizer and evaluated on standard benchmark datasets to compare accuracy before and after optimization. In cases where discrepancies were observed, we conducted pass-by-pass optimizations to isolate the specific transformations responsible for the faults.
@@
We focus on the ONNX Optimizer due to its impact and popularity (approximately $700$ GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier [5], which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. We aim to address this gap by introducing a differential testing methodology through OODTE. Our approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
@@
In total, we observed output differences between the original models and their and optimized counterparts in $34$ out of the $130$ models ($26.1$%). Additionally, we encountered issues with $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) optimizer passes - with either the optimizer crashing or generating an invalid model. Through detailed analysis, we identified $15$ distinct bugs, of which only one had been previously reported—and even then, only partially and without a conclusive explanation. In that case, we were able to provide substantial additional information not previously documented. All identified bugs were reported to the official ONNX Optimizer repository, accompanied by detailed descriptions, so that the optimizer developers can further examine and resolve them.
Overall, we present the following contributions:
1. OODTE  - a differential testing tool to examine the effects of the ONNX Optimizer to DNNs.
2. A survey of 130 models optimized using the ONNX Optimizer, presenting the effects of the optimizations on model correctness in detail.
3. A set of faults detected using OODTE - reported to the official ONNX Optimizer repository.
II Background
II-A Deep Learning Models
@@
II-A1 Image Classification
Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters. These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function [46].
II-A2 Object Detection and Semantic Segmentation
@@
II-A3 Text Comprehension and Generation
These models handle tasks such as summarization, recognition, and sentiment analysis. Text generation models, especially transformers [66], use contextual representations to iteratively predict the next token using a sliding window approach. This generative process continues until a stopping condition is met.
II-B Graph Optimizations
For deep neural networks (DNNs) to perform effectively on complex tasks, they must be optimized to operate more efficiently while maintaining their performance. A common approach to achieving this is optimizing the model graph to improve computational efficiency. This may involve techniques such as node pruning, node fusion, and constant folding, among others. Similar to conventional software, these optimizations are applied to DNNs, with AI compilers [12, 32] incorporating such techniques. Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
II-C The ONNX Optimizer
The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes. It consists of a set of $47$ passes grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency. Fusion passes merge operations (e.g., fuse_bn_into_conv, which merges BatchNorm and Conv nodes) to reduce overhead. Elimination passes remove redundant nodes, such as eliminate_deadend (removing nodes with unused outputs) and eliminate_nop_transpose (removing transpose nodes that do nothing). Graph rewrite passes modify graph structure for compatibility, such as rewrite_input_dtype, which adjusts input data types. By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
II-D Model Output Comparison Methods
@@
Classification models:
We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation [28], which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
Object Detection and Semantic Segmentation models:
Each prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score. For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity. For bounding boxes and confidence scores, we use standard metrics: (I) IoU for bounding box overlap, (II) F1 Score for overall detection balance, (III) Average Precision (AP) for precision across thresholds, (IV) Average Recall (AR) for recall across thresholds, (V) mAP for aggregate detection and localization accuracy. These metrics are computed at IoU thresholds of 0.5, 0.75, and 0.9, with the original model as the reference.
Text generation models:
We convert token outputs to plain text and use the BLEU Score [45] to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
Binary sentiment analysis models:
@@
III System Architecture & Methodology
@@
Figure 1: Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
@@
The architecture of OODTE with its different modules is shown in Figure 1. It has the following intuitive workflow: load models, optimize them, execute both the original and the optimized versions across the dataset and compare the results. If any discrepancies in model outputs are detected, OODTE  performs an iterative approach where each model under test is optimized with each ONNX Optimizer pass in isolation, and then the results are compared with the original model. This is done to localize the optimization pass responsible for causing an output difference. All results are stored in report files, both for reporting and localization purposes. In detail, OODTE  consists of the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. OODTE  loads model configuration settings in two ways: (1) via local configuration file, and (2) by fetching configuration via the ONNX repository from which the models are loaded from. This is managed by the Model Orchestrator Module. The modules within OODTE  in Figure 1 is explained in the following Sections.
III-A Model Orchestrator Module
@@
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models for optimization and execution. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading of an online repository. At the moment, OODTE supports the official ONNX Model Hub [7], however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
III-B Optimizer Module
@@
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations. This is useful across OODTE pipeline, when faulty behavior is discovered across model evaluation and the models presenting the faults need to be examined in depth. OODTE isolates such cases after a full pass optimization exeuction and performs per-pass optimization generation for that purpose.
Once a model is optimized with either setting, it then gets propagated to the Runner model, in order to perform inference.
III-C Runner Module
@@
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions. It then performs model inference, using ONNX Runtime [2], combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets [3] package.
@@
To enable the use of large-scale datasets and extract results from models that are larger and more computationally demanding—especially when operating on low-end devices—the Runner Module supports execution across a dataset in parts. In particular, a run can be divided into $N$ chunks, where each chunk consists of a fraction, $1/N$th, of the dataset. Any remaining data not evenly divisible is also included in the final chunk of the run. This process applies to both the original and the optimized model versions, with the results extracted and propagated to the Metrics Comparison Module for subsequent evaluation and reporting. By following this approach, we were able to conduct our survey on two consumer-grade devices without relying on a compute cluster to obtain results, as described in Section V.
III-D Metrics Comparison Module

### v3

@@ Abstract.
With over 760 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models. Despite its widespread use, its ability to maintain model accuracy during optimization has not been thoroughly investigated. In this work, we present OODTE, a utility designed to automatically and comprehensively evaluate the correctness of the ONNX Optimizer. OODTE adopts a straightforward yet powerful differential testing, fault localization, and evaluation methodology, which can be readily adapted for use with other compiler optimizers. Specifically, OODTE takes a collection of ONNX models, applies optimizations, and executes both the original and optimized versions across a user-defined input set, automatically capturing any issues encountered during optimization. When discrepancies in accuracy arise, OODTE iteratively isolates the responsible optimization pass by repeating the process at a finer granularity. We applied OODTE to 130 well-known models from the official ONNX Model Hub, spanning diverse tasks including classification, object detection, semantic segmentation, text summarization, question answering, and sentiment analysis. Our evaluation revealed that 9.29.2% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies. Additionally, 3030% of classification models and 16.616.6% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization. In total, OODTE uncovered 15 issues, 14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly. All findings were reported to the ONNX Optimizer development team. OODTE demonstrates a simple but effective framework for validating AI model optimizers and with minimal work, is applicable beyond the ONNX ecosystem.
1. Introduction
@@
Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models (speedisallyouneed, ; CompilerGym, ). Widely-used AI compilers and frameworks, such as Apache TVM (tvm, ) and MLIR (lattner2020mlircompilerinfrastructureend, ), incorporate their own native sets of optimization passes for efficient use of hardware. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
@@
Models can also be represented using industry-standard formats, such as ONNX (onnxsite, ). By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer (onnxoptimizer, ) serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question.
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. Ideally, such transformations should not introduce any deviations in model behavior. To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a suite that detects and localizes faults in the ONNX Optimizer.
Problem Being Solved
OODTE solves the problem of detection and localization of faults associated with the ONNX Optimizer, primarily in the direction of output label inconsistencies but also compiler crashes and miscompilations. In particular, OODTE employs a per-pass fault localization strategy combined with differential testing between original and optimized models of various types to identify and localize issues arising from specific passes in the ONNX Optimizer.
Case Studies
@@
To evaluate the effectiveness of OODTE, we collected 130 models as case studies, performing different tasks, such as classification, object detection, text comprehension, and text generation. All models were obtained from the official ONNX Model Hub (onnxmodelhub, ). Each model was optimized using the ONNX Optimizer and assessed on standard benchmark datasets to compare accuracy before and after optimization. When discrepancies were detected, we applied pass-by-pass optimization to isolate the specific transformations responsible for the faults.
@@
We focus on the ONNX Optimizer due to its impact and popularity (over 760 GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier (onnxsimplifier, ), which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. With OODTE, we aim to enhance the effectiveness of fault localization and detection in the optimizer. However approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
@@
By conducting experiments on our case studies, we observed output discrepancies between the original and optimized models in $34$ of $130$ cases (26.126.1%). We also encountered issues with $12$ model instances (9.29.2%) using the primary optimizer passes (fuse and eliminate) where the optimizer crashed or produced invalid models. Through analysis, we identified 15 distinct bugs, only one of which had been partially reported before (without a full explanation or reproduction steps). For this case, we provided additional insights. All bugs were reported to the official ONNX Optimizer repository with detailed descriptions for further resolution. OODTE effectively detected faults due to incorrect value references, input/output mismatches, and graph-structure inconsistencies introduced by the ONNX optimizer. These manifested as (1) optimizer crashes, (2) malformed models, and (3) valid models with output label inconsistencies compared to their unoptimized counterparts.
2. Background
2.1. Deep Learning Models
@@
2.1.1. Image Classification
Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters. These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function (softmax, ).
2.1.2. Object Detection and Semantic Segmentation
@@
2.1.3. Text Comprehension and Generation
These models handle tasks such as summarization, recognition, and sentiment analysis. Text generation models, especially transformers (attention, ), use contextual representations to iteratively predict the next token using a sliding window approach. This generative process continues until a stopping condition is met.
2.2. Graph Optimizations
For deep neural networks (DNNs) to perform effectively on complex tasks, they must be optimized to operate more efficiently while maintaining their performance. A common approach to achieving this is optimizing the model graph to improve computational efficiency. This may involve techniques such as node pruning, node fusion, and constant folding, among others. Similar to conventional software, these optimizations are applied to DNNs, with AI compilers (tvm, ; lattner2020mlircompilerinfrastructureend, ) incorporating such techniques. Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
2.3. The ONNX Optimizer
The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes. It consists of a set of $47$ passes grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency. Fusion passes merge operations (e.g., fuse_bn_into_conv, which merges BatchNorm and Conv nodes) to reduce overhead. Elimination passes remove redundant nodes, such as eliminate_deadend (removing nodes with unused outputs) and eliminate_nop_transpose (removing transpose nodes that do nothing). Graph rewrite passes modify the graph structure to ensure compatibility; for example, rewrite_input_dtype adjusts input data types. In this category, there are also secondary passes designed primarily for analysis and debugging, such as split_init and split_pre-
dict
, which isolate the initializer and inference subgraphs, respectively. By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
2.4. Model Output Comparison Methods
@@
Classification models::
We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation (kendall, ), which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
Object Detection & Semantic Segmentation models::
Ea-
ch prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score. For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity. For bounding boxes and confidence scores, we use standard metrics: (I) IoU for bounding box overlap, (II) F1 Score for overall detection balance, (III) Average Precision (AP) for precision across thresholds, (IV) Average Recall (AR) for recall across thresholds, (V) mAP for aggregate detection and localization accuracy. These metrics are computed at IoU thresholds of $0.5$, $0.75$, and $0.9$, with the original model as the reference.
Text generation models::
We convert token outputs to plain text and use the BLEU Score (BLEU, ) to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
Binary sentiment analysis models::
@@
3. System Architecture & Methodology
@@
Figure 1. Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
@@
The architecture of OODTE and its modules is shown in Figure 1. Its workflow is straightforward: load models, optimize them, execute both original and optimized versions across the dataset, and compare the results. If output discrepancies are detected, OODTE iteratively re-optimizes each model using individual ONNX Optimizer passes to localize the pass causing the difference. All results are stored in report files for analysis and fault localization.
OODTE consists of four primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. It loads model configuration settings either (1) from a local configuration file or (2) by fetching them from the ONNX repository from which the models are sourced, both managed by the Model Orchestrator Module. The modules in Figure 1 are detailed in the following sections.
3.1. Model Orchestrator Module
@@
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models for optimization and execution. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading models from an online repository. At the moment, OODTE supports the official ONNX Model Hub (onnxmodelhub, ), however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
3.2. Optimizer Module
@@
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model, applied in a user-defined order specified in the system configuration. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations. This feature is essential throughout the OODTE pipeline: when faulty behavior is discovered during model evaluation, the models exhibiting such faults must be examined in depth. To identify the pass responsible for problematic behavior, OODTE first compares fully optimized models with their originals throughout its pipeline, and upon detecting a crash or accuracy issue, isolates the problematic cases by generating per-pass optimizations and comparing them again with the originals in an iterative process, until the pass responsible for the issue is identified. Once a model is optimized with either setting, it then gets propagated to the Runner model for inference.
3.3. Runner Module
@@
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions. It then performs model inference, using ONNX Runtime (onnxruntime, ), combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets (datasets, ) package.
@@
To enable the use of large-scale datasets and extract results from models that are larger and more computationally demanding—especially when operating on low-end devices—the Runner Module supports execution across a dataset in parts. In particular, a run can be divided into $N$ chunks, where each chunk consists of a fraction, $1/N$th, of the dataset. Any remaining data not evenly divisible is also included in the final chunk of the run. This process applies to both the original and the optimized model versions, with the results extracted and propagated to the Metrics Comparison Module for subsequent evaluation and reporting. By following this approach, we were able to conduct our survey on two consumer-grade devices without relying on a compute cluster to obtain results, as described in Section 5.
3.4. Metrics Comparison Module